OpenStudy (anonymous):

What is the Indefinite integral of 1/(x^6sqrt(x^2-4)) using trigonometric substitution?

OpenStudy (anonymous):

\[\int\limits \frac{ dx }{ x^6\sqrt{x^2-4} }\]

OpenStudy (anonymous):

@anthonykanow put x=2sec theta

OpenStudy (anonymous):

yes. and \[\sqrt{x^2-a^2}=2\tan(\theta)\]

OpenStudy (anonymous):

\[\frac{ 2\sec(\theta)\tan(\theta) }{ 64\sec^6(\theta)2\tan(\theta) }\]= 1\[\frac{ 1 }{ 64 \sec^5(\theta) }\]

OpenStudy (anonymous):

then, \[\int\limits \frac{ 1 }{\sec^5(\theta)}dtheta\] yes?

OpenStudy (anonymous):

I mean \[1/64 \int\limits \frac{ 1 }{ \sec^5(\theta) } dtheta\]

OpenStudy (anonymous):

Yeah then convert sec to cos in the numerator

OpenStudy (anonymous):

@anthonykanow

OpenStudy (anonymous):

can you proceed further? it is very easy now.

OpenStudy (anonymous):

\[\int\limits_{}^{} \cos ^5 \theta d \theta \]

OpenStudy (anonymous):

hello...

OpenStudy (anonymous):

I know that 1/sec = cos I am having trouble with the integral of cos^5(theta)

OpenStudy (anonymous):

ok let me post a solution

OpenStudy (anonymous):

Ok, thank you

OpenStudy (michele_laino):

sorry I think a better substitution is: x=2 cosh(theta)

OpenStudy (anonymous):

excellent, I didn't think of that. Thank you

OpenStudy (anonymous):

@anthonykanow

OpenStudy (anonymous):

@Michele_Laino please post your solution, its good to do a problem in various ways

OpenStudy (michele_laino):

ok! I write

OpenStudy (michele_laino):

Sorry I'm not able to use the editor. I try to log out and then to log in

OpenStudy (anonymous):

do you have a math editor? then it will be very easy

OpenStudy (michele_laino):

here I am

OpenStudy (michele_laino):

\[x=2\cosh(\theta)\] so we can write: \[\frac{ 1 }{ 64 }\int\limits \frac{ d \theta }{ (senh(\theta))^{6} }\]

OpenStudy (anonymous):

next?

OpenStudy (michele_laino):

I'm trying...

OpenStudy (michele_laino):

next is: \[\frac{ 1 }{ 64 } \int\limits\limits \frac{ d \theta }{ (ch \theta) ^{6}}= \] \[=\frac{ 1 }{ 64 }\int\limits d(th(\theta))(1-th(\theta))^{4}\]

OpenStudy (michele_laino):

then after developing the fourth power, we can integrate easily

OpenStudy (michele_laino):

@Princer_Jones